Find the solution to the heat equation on the infinite domain

∂u∂t=k∂2u∂x2,−∞<x<∞,t>0,u(x,0)={x,0,|x|<1|x|>1.∂u∂t=k∂2u∂x2,−∞<x<∞,t>0,u(x,0)={x,|x|<10,|x|>1.

in terms of the error function.Q1 (10 points) Find the solution to the heat equation on the infinite domain azu ди at k -00<x<0, t>0, ar2 u(x,0) (X, 1x < 1
Answer #1

all 22 given heat equation 24 224 -K -0<x< 29 it>O at U(3,0) = sz, 18121 = f(x) 0, JX171 we know solution to this problem islet z = (x-8) rakt We know essor function erf (0) - 96 ex?dx dz = Jakt ? ds = 2- -J4 Ktdz Now z varies from Z= Xt1 Juka as 8-L kt -22 22 11 e d2 한d t AKE ez2d2 +) o H J4RA 본 JARt e22d2 t Jaktan E ē 22d2 제 해 Jakt NIRt 22 22 JÃ 01 - 현 ē dz 오 xt eff [esfrom equation K JqK 又 U(X,大) 示 erf J- / 十 eTV 2 (HK, 3 -(H) 體) 11 piecent colore *人 (4) *大 e (x) =(x+y)) UCXjk) - 宗 (行 「K 2 eAny doubt pls ask